OpenStudy (anonymous):

(sqrt)486 - (sqrt)24 + (sqrt) 6

OpenStudy (anonymous):

factor each integer, then pull any perfect squares outside of the radical

OpenStudy (anonymous):

for example \[486=2\times 3^5\] so \[\sqrt{486}=3^2\sqrt{2\times 3}=9\sqrt{6}\] each term will have a \[\sqrt{6}\] in it, then combine like terms

OpenStudy (anonymous):

is it 2(sqrt)6 or 8(sqrt)6